A motorcycle traveling at a speed of 36.0 mi/h needs a minimum of 41.0 ft to stop. If the same motorcycle is traveling 68.0 mi/h, determine its minimum stopping distance (in ft), assuming the same rate of acceleration.
Answer #1

here,

the initial speed of motorcycle , u1 = 36 mi/h

the initial minimum distance , s1 = 41 ft

when ,

the initial speed of motorcycle , u2 = 68 mi/h

the initial minimum distance is s2

final speed is zero

using third equation of motion

v^2 - u^2 = 2 * a * s

as v = 0

- u^2 = 2 * a * s

so,

s2/s1 = u2^2 /u1^2

s2 = 41 * 68^2 /36^2

s2 = 146.3 ft

the minimum distance is 146.3 ft